# Subwavelength arrays of quantum emitters: nonlinearities enter the weak-drive regime, and lead to correlated subradiant states

### v3

@@ Abstract
To solve the effective model, we use a method based on an expansion in the effective environment known as a non-crossing approximation (NCA), in a formulation that allows this to have both a Markovian and a non-Markovian component [61, 52, 62]. This approximation is controlled when the effect of neighbouring sites in the effective single-site problem is small, which eventually limits our approach to a maximum interaction strength (a minimum value of $k_{0}a$) above (below) which it does not converge. In the following we will set $k_{0}a=2$.
@@
In addition, we found that solving the DMFT equations with a simple fixed-point iteration scheme did not converge at weak drive intensities, for any $k_{0}a$ in the subwavelength regime. In this respect, we noticed that this scheme fails even in the simpler case of a Gutzwiller mean field theory, in the regime in which this predicts two simultaneous steady-states [45] – corresponding to a smaller $k_{0}a$ than considered here – as one of them corresponds to a repulsive fixed point. To circumvent these convergence problems, we adapted a linear mixing scheme and a Broyden algorithm estimating gradients [63], which can reach convergence at weaker drives. We discuss these methods more in detail in [59].
@@
Using DMFT/NCA, we predict a steady state with a finite density of subradiant excitation pairs. This is shown in Fig. 1 (c), showing the number of steady-state excitations resolved in momentum, $n_{c}(k)=\langle{\sigma_{k}^{+}\sigma_{k}^{-}}\rangle_{c}$, where $\sigma^{-}_{k}=\sum_{j}{\rm e}^{{\rm i}kr_{j}}\sigma^{-}_{j}/\sqrt{N}$, and for the driving scheme considered (Fig. 1 (b)) and at drive strengths of $\Omega\lesssim\Gamma$. Here the subscript “c” indicates a connected average $\langle{AB}\rangle_{c}\equiv\langle{AB}\rangle-\langle{A}\rangle\langle{B}\rangle$, which only affects $n_{c}(k)$ at $k=0$ and will be discussed later. Comparing with the overlaid dispersion relation, one sees that when the subradiant modes are resonant, for $\Delta\gtrsim V_{k_{0}}\approx 0.1\Gamma$, these are selectively populated. By contrast, when the radiative modes are resonant these do not get significantly populated as their finite lifetimes prevents the parametric instability to occur: instead, population concentrates at the boundary of the subradiant manifold, corresponding to the modes closest to resonance yet with the smallest decay rates. We remark that the subradiant steady-state population has a finite density, of order of $n_{\rm c,subrad}={\sum}_{{|k|>k_{0}}}\left\langle\sigma_{k}^{+}\sigma_{k}^{-}\right\rangle_{c}/N\approx 0.05$ for $\Omega/\Gamma\approx 1$ and $\Delta/\Gamma=0.4$, thus it is far from the zero-density single-excitation regime [25]. We also recall that such population is solely bound by non-linearities – as by the divergence of populations in the parametric instability of the weakly-non-linear model discussed – resulting in interactions among its constituent excitations. Yet, similarly as in the single-particle regime [25], this population has a well defined energy and momentum strongly-mismatched from free photons, thus it is subradiant.
@@
Figure 2: Pair correlation function of the steady state computed in DMFT/NCA, for the non-linear driving conditions illustrated in Fig. 1 and with a drive detuning of $\Delta/\Gamma=0.4$. The subscript “c”(“d”) indicates the connected (disconnected) component of such function, describing its non-linear (linear) component. (a) The connected function as a function of momentum $k$ for different drive strengths $\Omega$: it shows two-mode squeezing correlations of mode-pairs with momenta (k,k)(-k,k). Decreasing the drive strength it becomes more sharply peaked around the resonant modes. (b) The non-local component of the connected function in real space, where iji-j is the lattice-sites distance, showing long-range correlations. (a) and (b) also show that correlations increase in magnitude as the drive is decreased, as a non-trivial interplay of reduced occupations that bound correlations, and reduced heating. (c) The ratio $W_{\rm subrad}$ (with a minus sign) between the connected component summed over the subradiant modes $c_{c,\rm subrad}=\underset{{|k|>k_{0}}}{\sum}\left\langle\sigma_{k}^{-}\sigma_{-k}^{-}\right\rangle_{c}/N$ and the disconnected one $c_{\rm d}=\left\langle\sigma_{k=0}^{-}\right\rangle\left\langle\sigma_{k=0}^{-}\right\rangle/N$, showing that the non-linear contribution to the steady-state is far from a small perturbation of the linear one.
@@
In addition to populating the subradiant manifold, the non-linear processes involved create excitations in entangled pairs. Nevertheless, such correlations could in principle be suppressed by heating due to interactions, and by the coupling with the environment. To assess this, we computed the connected pair correlation function $\langle{\sigma^{-}_{k}\sigma^{-}_{-k}}\rangle_{c}$. Such function and its real-space counterpart $\langle{\sigma^{-}_{i}\sigma^{-}_{j}}\rangle_{c}$ are shown in panels (a) and (b) of Fig. 2. Panel (a) shows that correlations are indeed present in mode-pairs with momenta k-k and $k$, peaking around the resonant modes for decreasing drive strengths. These correspond to correlations that are long-ranged in real space, as shown in panel (b). Both panels also show a second important fact: the correlations amplitude increases decreasing the drive strength. This results from a non-trivial interplay between decreased populations and decreased heating. It is then clear that correlations can survive heating effects due to drive, dissipation and interactions in such weak-drive steady-state. Decreasing the drive further, eventually a maximum will be reached and the trend will invert, as correlations become bound by occupations. For those smaller drive strengths though, our method is not reliable and predicts negative probabilities due to an expected breakdown of the non-crossing approximation (NCA). For this reason, we limited our analysis to intermediate drive intensities for which our approach is well behaved. We also remark that, although these correlations are reminiscent of a two-mode bosonic squeezed state, the steady-state predicted is non-gaussian. Finally, we note that a qualitatively similar behaviour is also displayed by the correlator $\langle{\sigma^{+}_{k}\sigma^{-}_{k}}\rangle_{c}$.
@@
Moreover, we wish to quantify the strength of non-linearities in the weak-drive regime. In particular, we consider their effect on the steady-state. This in fact features also an uncorrelated, linear component. This is due to the linear driving, although detuned, of the zero-momentum mode in our geometry, which results into a coherent population of this mode witnessed by a non-zero, uniform $\langle{\sigma_{j}^{-}}\rangle$. It is the only component that would be predicted by classical theories [40, 41].
@@
We estimate the relative importance of the non-linear and linear components taking the ratio $W(A)=-c_{\rm c,A}/c_{\rm d}$, between the connected pair-correlation function summed over a momentum region $A$, $c_{\rm c,A}=\sum_{k\in A}\langle{\sigma^{-}_{k}\sigma^{-}_{-k}}\rangle_{c}/N$, and the disconnected component, $c_{\rm d}=\left\langle\sigma_{k=0}^{-}\right\rangle\left\langle\sigma_{k=0}^{-}\right\rangle/N$, carried by a delta function contribution at $k=0$. We note that if $A$ is the entire Brillouin zone, then the numerator is rigidly related to the denominator by a sum rule $c_{\rm c}=\langle{\sigma^{-}_{j}}\rangle\langle{\sigma^{-}_{j}}\rangle=c_{d}$, and $W=1$ regardless of the state. This results from the property $\langle{\sigma^{-}_{i}\sigma^{-}_{i}}\rangle=0$ and is respected in DMFT/NCA. This sum rule shows that $W$ defines a distribution of the non-linear correlations in momentum space that is normalised, such that it depends on the drive strength but its overall amplitude doesn’t, and is thus interesting to track as a function of drive strength. In particular, we define $W_{\rm subrad}$ taking $A$ equal to the subradiant manifold ($A=\{k:\left\lvert k\right\rvert>k_{0}\}$), which compares its non-linear weight with the linear component and which is plotted in Fig. 2 (c). This figure shows that $W_{\rm subrad}$ is about $1/2$, indicating a strongly non linear state. Also, the trend with drive strength shows that the state becomes the more non-linear the weaker the drive.
@@
Although here we focused on the steady state, we also remark that the linear uncorrelated component of the steady-state is superradiant and would be quickly dissipated by switching the drive off, leaving behind a slowly-decaying correlated subradiant state. In the Supplemental Material [59] we also discuss an important role played by non-linearities in the relaxation of dynamical correlation functions, rather than in the steady-state density matrix.
Discussion – Our results demonstrate a novel, unexpected nonlinear regime of quantum emitters ensembles, a long-standing quantum many-body problem in atomic physics [1, 2, 3, 4]. In fact, the weak-drive regime of these systems had been widely regarded as a linear regime, well described by classical equations of motion [40, 41, 42, 43, 44, 12, 45, 46, 47, 48, 49]. We demonstrated that this is instead strongly non-linear in the case of large subwavelength arrays, due to the important contribution of resonant higher-order processes exciting single-particle subradiant mode-pairs. Using a Dynamical Mean-Field Theory (DMFT), we predicted that this results into a steady-state of interacting, subradiant excitations pairs, characterised by correlations with two-mode squeezing and long-range character.
@@
Our predictions can be realised in different experimental platforms. Namely, they can be tested in current experiments with semiconductor dipolar excitons [9], which already operate in suitable subwavelength regimes. They could also be implemented in upcoming experiments with ultracold atoms [8, 7] that are close to achieving suitable lattice spacings. Both platforms directly implement the model studied.
Our results have significant potential for applications. They demonstrate that the weak-drive regime of subwavelength quantum emitter arrays ought to be reconsidered with renewed interest. This is an especially interesting regime, as it combines strong non-linearities with weak driving powers. As a result, it opens new avenues for non-linear quantum optics [18, 19, 17, 20]. For example, it could be used for the generation of entangled photon pairs: in finite-size chains, the correlated pairs of subradiant atomic excitation, created via the non-linear mechanism described, can scatter into outgoing entangled photons at the chain boundaries – where subradiant modes couple to propagating photon modes [25].
In addition, the combination of weak driving and strong non-linearities allows to prepare quantum-correlated states of the emitters, and possibly to control them coherently, while avoiding significant heating effects that would occur in more conventional nonlinear regimes at strong driving. This lays the ground for exploring, controlling and harnessing such correlations. In particular, the predicted steady-state, featuring multi-mode squeezing correlations together with a simple preparation scheme, has significant potential for quantum metrology. There, states with this kind of correlations are highly sought after [53], but finding viable preparation schemes is a major challenge [64].
Our predictions also reveal a simple, experimentally viable route to populate subradiant states of quantum emitter arrays, which does not require additional single-atom resolution [39, 50, 29, 32, 33, 34] which is difficult to achieve in subwavelength regimes. By contrast, the non-linear mechanism here described only requires the application of a far-field drive. This contributes to enabling sought after applications of subradiant modes, including for photon storage [28, 29, 25, 30, 31, 32, 33, 34] and for quantum metrology [35, 36, 37].
Finally, we remark that, at a technical level, obtaining consistent predictions for large arrays in the weak-drive regime in which subradiant states play a significant role is a remarkable achievement, as this quest had defeated previous theoretical methods [65, 66, 27]. It was possible thanks to the steady-state DMFT approach deployed [51, 52], and further improved in the present work – although with the limitations discussed.
@@
Supplemental Material for:
“Subwavelength arrays of quantum emitters:
nonlinearities enter the weak-drive regime, and lead to correlated subradiant states”
This Supplemental Material is organized as follows. In Sec. A we take non-linearities into account perturbatively, and show that this leads to an instability of the steady state. In Sec. B we discuss the main equations of the DMFT approach, in Sec. C the numerical schemes to solve them, and in Sec. D the exact limits of our DMFT approach. In Sec. E we discuss the linear stability of the DMFT steady-state solution. Finally, in Sec. F we discuss the effect of interactions on time-dependent correlation functions, rather than on the steady-state density matrix.
Appendix A Weakly-nonlinear theory of the weak-drive regime
Here we develop a weakly non-linear theory for the steady-state, in which non-linearities are taken into account perturbatively with respect to the non-interacting theory, assuming a weak drive strength $\Omega$ and a large array size $N$. This is similar in some aspects to the treatment of a weakly-interacting Bose gas [67].
@@ Appendix A Weakly-nonlinear theory of the weak-drive regime
$\begin{split}H=&\sum_{k}(-\Delta+V_{k})a^{\dagger}_{k}a_{k}+\frac{\sqrt{N}\Omega}{2}(a_{k=0}+a_{k=0}^{\dagger})\\ &-\sum_{kq}\frac{\Omega}{4\sqrt{N}}(a_{k-q}^{\dagger}a_{q}^{\dagger}a_{k}+\text{h.c.})\\ &-\sum_{kqp}\frac{V_{k}}{2N}(a_{q-k-p}^{\dagger}a_{p}^{\dagger}a_{q}a_{k}+\text{h.c.}).\end{split}$ (7)
Regarding both drive and non-linear terms as perturbations in a time-dependent perturbation theory, populating the modes with $V_{\pm{k}_{\rm res}}\neq\Delta$ involves non-resonant processes that are energetically suppressed. We therefore retain only the resonant modes $\pm k_{\rm res}$ – and drop the ”res” label hereafter – such that the problem reduces to a three-modes problem (including the driven $k=0$ mode).
We then incorporate the non-linearities perturbatively with respect to the non-interacting steady-state under coherent driving. In absence of non-linearities, the steady-state for the $k=0$ mode is a coherent state with
@@
which is proportional to $\sqrt{N}$ due to the direct coupling with the drive. For simplicity, we further assume $\sqrt{N}\Omega\ll\Gamma_{k=0}\sim O(1)$ such that powers of $a_{k=0}$ larger than two can be neglected. Then, the fluctuations of $a_{k=0}$ in the steady-state are only determined by its coupling with the other modes through the non-linearities, as these vanish for the coherent non-interacting steady state. Provided that these are negligible with respect to its macroscopic average value $\sim\Omega\sqrt{N}$, we can replace the operator $a_{k=0}$ with its average. A posteriori, the non-interacting Hamiltonian of the $k=0$ mode is of order $N\Omega^{2}$, and the non-linear terms $\frac{\Omega}{4\sqrt{N}}a_{0}a^{\dagger}_{k}a^{\dagger}_{-k}$ and $\frac{V_{k}}{2N}a_{0}a_{0}a_{k}^{\dagger}a_{-k}^{\dagger}$ that would lead to fluctuations of $a_{0}$ can be neglected if $a_{\pm k}\lesssim\sqrt{N}$. The substitution then leads to the weakly-interacting Hamiltonian
@@
$H_{k}$ $\approx\left(a_{k}^{\dagger}a_{k}+a_{-k}^{\dagger}a_{-k}\right)\left[\left(\text{Re}^{2}\tilde{\alpha}_{0}+2|\tilde{\alpha}_{0}|^{2}V_{0}\right)\right]$ (9)
$+\left(a_{k}^{\dagger}a_{-k}+a_{-k}^{\dagger}a_{k}\right)\left(-2\Omega^{2}|\tilde{\alpha}_{0}|^{2}V_{k}\right)$
@@
$+\left[\left(a_{k}^{\dagger}a_{k}^{\dagger}+a_{-k}^{\dagger}a_{-k}^{\dagger}\right)\left(-\Omega^{2}\frac{V_{k}}{2}\tilde{\alpha}_{0}^{2}\right)+\text{h.c.}\right]$
$+\left[\left(a_{-k}^{\dagger}a_{k}^{\dagger}+a_{k}^{\dagger}a_{-k}^{\dagger}\right)\left(-\frac{\Omega^{2}}{4}\tilde{\alpha}_{0}^{2}-\Omega^{2}\frac{V_{0}}{2}\tilde{\alpha}_{0}^{2}\right)+\text{h.c.}\right]$
We notice that all contributions are of leading order $\Omega^{2}$, and independent of $N$.
@@
The same kind of treatment is then applied to the dissipator. For a drive resonant with the subradiant modes, the decay rate of these modes is already small for large arrays – $\Gamma\sim N^{-\alpha}$ ($\Gamma\sim\exp(-N)$) for open (closed) boundary conditions –, thus one can simply replace $\sigma_{k}^{-}\rightarrow a_{k}$ in Eq. (3), neglecting further terms from the Holstein-Primakoff expansion.

### v4

@@ Abstract
To solve the effective model, we use a method based on an expansion in the effective environment and a resummation of non-crossing diagrams (NCA), in a formulation that allows both a Markovian and a non-Markovian component [57, 54, 56]. This is controlled when the effect of neighbouring sites in the effective single-site problem is small, which eventually limits our approach to a maximum interaction strength (a minimum value of $k_{0}a$) above (below) which it does not converge. In the following we will set $k_{0}a=2$.
@@
We found that solving the DMFT equations with a simple fixed-point iteration scheme did not converge at weak drive intensities, for any $k_{0}a$ in the subwavelength regime. In this respect, this scheme fails even in the simpler case of a Gutzwiller mean-field theory, in the regime in which this predicts two simultaneous steady-states [45] – though this corresponds to smaller $k_{0}a$s than considered here –, as one of them corresponds to a repulsive fixed point. To circumvent similar convergence problems, we implemented gradient-based methods (and/or a linear mixing scheme), which can in principle reach convergence at all fixed points (see SM Note C).
@@
Using DMFT/NCA, we predict a steady state with a finite density of subradiant excitation pairs. This is shown in Fig. 1 (c), reporting the number of steady-state excitations resolved in momentum, $n_{c}(k)=\langle{\sigma_{k}^{+}\sigma_{k}^{-}}\rangle_{c}$, where $\sigma^{-}_{k}=\sum_{j}{\rm e}^{{\rm i}kr_{j}}\sigma^{-}_{j}/\sqrt{N}$, and for the driving scheme considered (Fig. 1 (b)) and at drive strengths of $\Omega\lesssim\Gamma$. Here the subscript “c” indicates a connected average $\langle{AB}\rangle_{c}\equiv\langle{AB}\rangle-\langle{A}\rangle\langle{B}\rangle$, which only affects $n_{c}(k)$ at $k=0$ and will be discussed later. Comparing with the overlaid dispersion relation, one sees that when the subradiant modes are resonant – for $\Delta\gtrsim V_{k_{0}}\approx 0.1\Gamma$ – these are selectively populated. By contrast, when the radiative modes are resonant these do not get significantly populated as their finite lifetimes prevents the parametric instability to occur: instead, the population concentrates at the boundary of the subradiant manifold, corresponding to the modes closest to resonance yet with the smallest decay rates. The linear stability of the steady-state was checked by enforcing a criterion for non-homogeneous fluctuations to remain bounded (see SM Note E). We remark that the subradiant steady-state population has a finite density, of order of $n_{\rm c,subrad}={\sum}_{{|k|>k_{0}}}\left\langle\sigma_{k}^{+}\sigma_{k}^{-}\right\rangle_{c}/N\approx 0.05$ for $\Omega/\Gamma\approx 1$ and $\Delta/\Gamma=0.4$, thus it is far from the zero-density, single-excitation regime [4]. We also recall that such population is solely bound by nonlinearities – as it diverges in the effective linear model obtained perturbatively – resulting in interactions among its constituent excitations. Yet, similarly as in the single-particle regime [4], this population has a well defined energy and momentum strongly-mismatched from free photons, thus it is subradiant.
@@
Figure 2: Pair-correlations function of the steady state computed in DMFT/NCA, for the driving conditions illustrated in Fig. 1 and with a drive detuning of $\Delta/\Gamma=0.4$. The subscript “c”(“d”) indicates the connected (disconnected) component of such function, describing its nonlinear (linear) component. (a) The connected function as a function of momentum $k$ for different drive strengths $\Omega$: it shows multimode squeezing correlations of mode-pairs with momenta (k,k)(-k,k). Decreasing the drive strength it becomes more sharply peaked around the resonant modes. (b) The non-local component of the connected function in real space, where iji-j is the lattice-sites distance, showing long-range correlations. (a) and (b) also show that correlations increase in magnitude as the drive is decreased, as a non-trivial interplay of reduced occupations – bounding correlations –, and reduced heating. (c) The ratio $W_{\rm subrad}$ (with a minus sign) between the connected component summed over the subradiant modes $c_{c,\rm subrad}=\underset{{|k|>k_{0}}}{\sum}\left\langle\sigma_{k}^{-}\sigma_{-k}^{-}\right\rangle_{c}/N$ and the disconnected one $c_{\rm d}=\left\langle\sigma_{k=0}^{-}\right\rangle\left\langle\sigma_{k=0}^{-}\right\rangle/N$, showing that the nonlinear contribution to the steady-state is far from a small perturbation.
@@
In addition to populating the subradiant manifold, the nonlinear processes involved create excitations in correlated pairs. Nevertheless, such correlations could in principle be suppressed by heating due to interactions, driving and coupling with the vacuum environment. To assess this, we computed the connected pair-correlations function $\langle{\sigma^{-}_{k}\sigma^{-}_{-k}}\rangle_{c}$. Such function and its real-space counterpart $\langle{\sigma^{-}_{i}\sigma^{-}_{j}}\rangle_{c}$ are shown in panels (a) and (b) of Fig. 2. Panel (a) shows that correlations are indeed present in mode-pairs with momenta k-k and $k$, peaking around the resonant modes for decreasing drive strengths. These correspond to correlations that are long-ranged in real space, as shown in panel (b). Both panels also show a second important fact: the correlations amplitude increases decreasing the drive strength. This results from a non-trivial interplay between decreased populations and decreased heating. It is then clear that correlations can survive heating effects in such weak-drive regime. Decreasing the drive further, we expect eventually a maximum and a trend inversion, as correlations become bound by occupations. For those smaller drive strengths though, our method is not reliable and predicts negative probabilities due to an expected breakdown of the non-crossing approximation (NCA). For this reason, we limited our analysis to intermediate drive intensities for which our approach is well behaved. We also remark that, although these correlations are reminiscent of a multimode bosonic squeezed state, the steady-state predicted is non-gaussian. A qualitatively similar behaviour was also found for the correlator $\langle{\sigma^{+}_{k}\sigma^{-}_{k}}\rangle_{c}$.
@@
Moreover, we wish to quantify the strength of nonlinearities in the weak-drive regime, and we consider their effect on the steady-state. This state in fact features also an uncorrelated linear component, due to the linear driving, although detuned, of the zero-momentum mode, resulting in a coherent population witnessed by a uniform $\langle{\sigma_{j}^{-}}\rangle\neq 0$. It is the only component that would be predicted by classical theories [53, 29].
@@
We estimate the relative importance of the nonlinear and linear components taking the ratio $W(A)=-c_{\rm c,A}/c_{\rm d}$, between the connected pair-correlation function summed over a momentum region $A$, $c_{\rm c,A}=\sum_{k\in A}\langle{\sigma^{-}_{k}\sigma^{-}_{-k}}\rangle_{c}/N$, and the disconnected component, $c_{\rm d}=\left\langle\sigma_{k=0}^{-}\right\rangle\left\langle\sigma_{k=0}^{-}\right\rangle/N$, carried by a delta function contribution at $k=0$. We note that if $A$ is the entire Brillouin zone, then the numerator is rigidly related to the denominator by a sum rule $c_{\rm c}=\langle{\sigma^{-}_{j}}\rangle\langle{\sigma^{-}_{j}}\rangle=c_{d}$, and $W=1$ regardless of the state. This results from the property $\langle{\sigma^{-}_{i}\sigma^{-}_{i}}\rangle=0$ and is respected in DMFT/NCA. This sum rule shows that $W$ defines a distribution of the nonlinear correlations in momentum space that is normalised, such that it depends on the drive strength but its overall amplitude doesn’t, and is thus interesting to track as a function of drive strength. In particular, we define $W_{\rm subrad}$ taking $A$ equal to the subradiant manifold ($A=\{k:\left\lvert k\right\rvert>k_{0}\}$), which compares its nonlinear weight with the linear component and which is plotted in Fig. 2 (c). This figure shows that $W_{\rm subrad}\approx 1/2$, indicating a strongly nonlinear state. This indicates a high conversion efficiency of drive photons into the nonlinearly populated modes relevant for nonlinear optics. In addition, the trend with drive strength shows that the state becomes the more nonlinear the weaker the drive.
@@
Although here we focused on the steady state, we also remark that the linear, uncorrelated component of the steady-state is superradiant and would be quickly dissipated by switching the drive off, leaving behind a slowly-decaying correlated subradiant state. In addition, nonlinearities also play and important role in shaping the dynamical correlation functions – rather than the steady-state density matrix. This is shown in SM Note F, where the two-point local correlations are shown to fail to relax altogether if nonlinearities are neglected, thus those in turn completely determine the relaxation behaviour.
Conclusions – We demonstrated that large subwavelength arrays of quantum emitters display a non-perturbative and strongly-nonlinear response down to arbitrary-weak drive intensities. This is surprising, because the weak-drive regime of these systems had been widely regarded as a linear one, well described by classical equations [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26]. This effect is due to the important contribution of higher-order, resonant processes involving pairs of single-particle subradiant modes, which has been overlooked in previous analyses. It depicts a scenario in stark contrast with the typical one of nonlinear optics, in which thick media and large drive powers are needed.
Using a Dynamical Mean-Field Theory (DMFT), we showed that such regime leads to a steady state of interacting pairs of subradiant excitations, characterised by correlations with multimode squeezing and long-range character.
At a technical level, obtaining consistent predictions at weak driving is notable, as this has proven extremely challenging for theoretical methods [42, 32, 50]. Our DMFT/NCA approach was successful, albeit with the limitations discussed.
@@
Our predictions can be realised in multiple experimental platforms, which directly implement the studied model: semiconductor dipolar excitons [35], already probing deep-subwavelength physics, and ultracold atoms [52, 62], which will soon operate in suitable regimes.
Our results have significant potential for applications. First, they reveal a simple, experimentally-viable route to populate the subradiant states of quantum emitter arrays. These are relevant for several purposes, notably photon storage [46, 17, 4, 43, 19, 6, 51, 13] and quantum metrology [44, 18, 48]. However, their excitation is challenging, motivating a range of proposals [67, 30, 17, 6, 51, 13] that nevertheless require fine experimental control such as single-atom resolution [67, 30, 17, 6, 51, 13], which is difficult to achieve in subwavelength regimes—in addition to the proposals being restricted to single-particle manipulation. By contrast, the nonlinear mechanism here discovered only requires a far-field coherent drive, routinely available in experiments.
Second, our results demonstrate that the weak-drive regime of subwavelength quantum emitter arrays is far from being the classical regime previously thought [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26], and it needs to be reconsidered. This is particularly relevant as it combines strong nonlinearities with weak driving powers and atom-thin structures, thereby opening a new avenue for nonlinear quantum optics [33, 39, 61, 65]: by avoiding the significant heating inherent to conventional strong-driving regimes, this approach enables substantially stronger quantum correlations—as evidenced by the predicted correlated steady state—and supports far richer nonlinear quantum behaviour.
In this respect, the one-dimensional setup considered here may enable new avenues for entangled-photon generation [33, 39, 61, 65]: the correlated pairs of subradiant atomic excitations can scatter into outgoing photons at the boundaries of a finite chain, where subradiant modes—despite being overall dark—couple to propagating photons [4]. The outgoing photons can thereby be collected, potentially inheriting intricate multimode correlations from the atomic excitations.
Finally, our prediction of a steady state exhibiting multimode squeezing correlations, combined with a clear and simple preparation scheme, is directly relevant to quantum metrology. Although such correlations are highly sought after [40], viable preparation strategies remain scarce [34].
@@
Supplementary Material for:
“How Subradiance Enables Nonlinearity in Weakly Driven Quantum Arrays”
Orazio Scarlatella1,2 and Nigel R. Cooper1
1TCM Group, Cavendish Laboratory, University of Cambridge, Cambridge CB3 0HE, UK
2Sorbonne Université, CNRS, Institut des NanoSciences de Paris, 75005 Paris, France
This Supplementary Material is organized as follows. In SM Note A we take nonlinearities into account perturbatively, and show that this leads to an instability of the steady state. In SM Note B we discuss the main equations of the DMFT approach, in SM Note C the numerical schemes to solve them, and in SM Note D the exact limits of our DMFT approach. In SM Note E we discuss the linear stability of the DMFT steady-state solution. Finally, in SM Note F we discuss the effect of interactions on time-dependent correlation functions, rather than on the steady-state density matrix.
SM Note A Weakly-nonlinear theory of the weak-drive regime
Here we develop a weakly nonlinear theory for the steady-state, in which nonlinearities are taken into account perturbatively with respect to the noninteracting theory, assuming a weak drive strength $\Omega$ and a large array size $N$. This is to some extent similar to the treatment of a weakly-interacting Bose gas [12].
@@
$\begin{split}H=&\sum_{k}(-\Delta+V_{k})a^{\dagger}_{k}a_{k}+\frac{\sqrt{N}\Omega}{2}(a_{k=0}+a_{k=0}^{\dagger})\\ &-\sum_{kq}\frac{\Omega}{4\sqrt{N}}(a_{k-q}^{\dagger}a_{q}^{\dagger}a_{k}+\text{h.c.})\\ &-\sum_{kqp}\frac{V_{k}}{2N}(a_{k+q-p}^{\dagger}a_{p}^{\dagger}a_{q}a_{k}+\text{h.c.}).\end{split}$ (7)
Regarding both drive and nonlinear terms as perturbations in a time-dependent perturbation theory, populating the modes with $V_{\pm{k}}\neq\Delta$ involves non-resonant (i.e. energy non-conserving) processes that are energetically suppressed. We therefore retain only the resonant modes with $k=k_{\rm res}$ – and drop the “res” label hereafter – such that the problem reduces to a three-modes problem (including the driven $k=0$ mode).
We then use this model as the basis for a perturbative analysis, with respect to the noninteracting model discussed in the main text – the first line of (7). First, we notice that the fluctuations of the $k=0$ mode around its average can be safely neglected. Indeed, in absence of nonlinearities, the steady-state for the $k=0$ mode is a coherent state with
@@
which is proportional to $\sqrt{N}$ due to the direct coupling with the drive. We further assume $\sqrt{N}\Omega\ll\Gamma_{k=0}$ such that $a_{k=0}$ is small, and terms containing powers larger than two in this operator can be neglected. Then, the fluctuations of $a_{k=0}$ in the steady-state are only determined by its (nonlinear) coupling with the other modes – as they vanish in the coherent noninteracting steady state. Provided that these fluctuations are negligible with respect to the (macroscopic) average value $\alpha_{0}\sim\Omega\sqrt{N}$ we can replace the operator $a_{k=0}$ with its average. A posteriori, the nonlinear terms $\frac{\Omega}{4\sqrt{N}}a_{0}a^{\dagger}_{k}a^{\dagger}_{-k}$ and $\frac{V_{k}}{2N}a_{0}a_{0}a_{k}^{\dagger}a_{-k}^{\dagger}$, which would lead to fluctuations of $a_{0}$, can also be neglected if $a_{\pm k}\lesssim\sqrt{N}$, as this ensures that they are small compared to the noninteracting Hamiltonian of the $k=0$ mode, of order $N\Omega^{2}$ – this drive-independent threshold can always be enforced at a sufficiently weak driving, as $a_{\pm k}$ must decrease with the drive strength. The substitution $a_{k=0}\rightarrow\alpha_{0}$ then leads to an effective nonlinear model for the $a_{\pm k}$ modes.
To proceed with a perturbative analysis, one must be able to treat the further nonlinear terms in the operators $a_{\pm k}$ perturbatively, with respect to a noninteracting component of the model. Such a noninteracting part here is
@@
$H_{k}$ $\approx-\left(a_{k}^{\dagger}a_{k}+a_{-k}^{\dagger}a_{-k}\right)\Omega^{2}\left(\text{Re}\tilde{\alpha}_{0}+2|\tilde{\alpha}_{0}|^{2}V_{0}+2|\tilde{\alpha}_{0}|^{2}V_{k}\right)$ (9)
@@
$-\left[a_{-k}^{\dagger}a_{k}^{\dagger}\;\Omega^{2}\tilde{\alpha}_{0}^{2}\left(\frac{1}{2}+{V_{0}}+{V_{k}}\right)+\text{h.c.}\right],$
@@
where all contributions are of leading order $\Omega^{2}$, and independent of $N$. The same treatment is applied to obtain the dissipator. For a drive resonant with the subradiant modes, the decay rate of these modes is already small for large arrays – $\Gamma\sim N^{-\alpha}$ ($\Gamma\sim\exp(-N)$) for open (closed) boundary conditions –, thus one can simply replace $\sigma_{k}^{-}\rightarrow a_{k}$ in Eq. (3), neglecting further terms from the Holstein-Primakoff expansion.